Calculate the heat of reaction in (kJ) for the following reaction:

2CO(g) + O2(g) --> 2CO2(g)

given the following data:

C + 1/2O2 --> CO(g) delta H= -110.5 KJ/mol

C + O2(g) --> CO2(g) delta H= -393.5 KJ/ mol
Answer #1

Reaction 1 : C (s) + 1/2 O2 (g) CO (g) : H1 = -110.5 kJ/mol

Reaction 2 : C (s) + O2 (g) CO2 (g) : H2 = -393.5 kJ/mol

Desired reaction : 2 CO (g) + O2 (g) 2 CO2 (g)

Step 1 : Reverse Reaction 1

Reaction 3 : CO (g) C (s) + 1/2 O2 (g)

H3 = -(H1)

H3 = -(-110.5 kJ/mol)

H3 = 110.5 kJ/mol

Step 2 : Add Reaction 2 and Reaction 3

Reaction 4 : CO (g) + 1/2 O2 (g) CO2 (g)

H4 = (H2) + (H3)

H4 = (-393.5 kJ/mol) + (110.5 kJ/mol)

H4 = -283.0 kJ/mol

Step 3 : Multiply Reaction 4 by a factor of 2

Reaction 5 : 2 CO (g) + O2 (g) 2 CO2 (g)

This is the desired reaction

H5 = 2 * (H4)

H5 = 2 * (-283.0 kJ/mol)

H5 = -566.0 kJ/mol

Heat of reaction is -566.0 kJ